H2S, a toxic gas with rotten egg like smell is used for qualitative analysis. If the
solubility of H2S in water at STP is 0.195 m, calculate Henry’s law constant

Grade:upto college level

1 Answers

Caleb Mamidi
6 years ago
pressure of h2s=kh*molefraction of h2s
molefraction of h2s,x=0.195/0.195+55.56=0.0035
kh=p/x =0.987/0.0035 = 282bars